OpenStudy (anonymous):

Find the domain of f for each of the following functions. If you need to enter - or , enter -INFINITY or INFINITY. f(x) = x+4 ------ x^3-9x a. (-INFINITY,0)(3,INFINITY) b. ALL REAL NUMBERS c. ALL REAL NUMBERS - {3} d. ALL REAL NUMBERS - {-3,3,0}

OpenStudy (dumbcow):

x^3 -9x = 0 x(x^2 -9) =0 x(x+3)(x-3) =0 D

OpenStudy (anonymous):

so you always set the denominator to 0 & do nothing with the numerator?